OpenStudy (anonymous):

abs val of i*z(conjugate) is equal to 2sqrt2 z=a+i i^2=-1 a=?

OpenStudy (anonymous):

i am not getting this problem. is it \[z=a+i, \overline z = a-i, |i^{\overline z}|=2\sqrt{2}\]?

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

sorry third

OpenStudy (anonymous):

i wrote the whole problem. is that the whole thing?

OpenStudy (anonymous):

that is, you are sure that \[z=a+i\] and so \[\overline z =a-i\] rigtht

OpenStudy (anonymous):

yes its the whole thing

OpenStudy (anonymous):

then i am confused. you have \[i=e^{\frac{\pi}{2}}\] \[i^{a-i}=e^{\frac{\pi}{2}i(a-i)}\] \[=e^{\frac{\pi}{2}}e^{\frac{a\pi}{2}i}\] and it looks to me like the absolute value of this thing is \[e^{\frac{\pi}{2}}\]so maybe i am missing something

OpenStudy (anonymous):

typo on first line should read \[i=e^{\frac{\pi}{2}i}\]

OpenStudy (anonymous):

hold on, is this \[i\times \overline z\] or \[i^{\overline z}\]?

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

if it is the first one then no problem. yoju get \[i(a-i)=1+ai\] so \[|1+ai|=\sqrt{1^2+a^2}=2\sqrt{2}\] and so \[1+a^2=8\] making \[a^2=7\] and thus \[a=\pm\sqrt{7}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw